Inventory information for Part 311 of Bonita Corp. discloses the following information for the month of June.

June   1

Balance

304 units @ $13

June 10

Sold

205 units @ $30

11

Purchased

804 units @ $15

15

Sold

504 units @ $32

20

Purchased

498 units @ $16

27

Sold

295 units @ $34

Cost of goods sold

(a) Periodic system, FIFO cost flow $ (b) Perpetual system, FIFO cost flow $ (c) Periodic system, LIFO cost flow $ (d) Perpetual system, LIFO cost flow $ (e) Periodic system, weighted-average cost flow $ (f) Perpetual system, moving-average cost flow $
Answer #1

10 un Purchase Sales Date Units Price Total cost Date Units 1-Jun 304 13 3952 10-Jun 205 11-Jun 804 15 12060 15-Jun 504 20-JuPerpetual method FIFO Method Closing inventory = ( 498 * 16 ) + ( 104 * 15 ) ======= Cost of goods sold = 23980 - 9528 ======